OpenStudy (anonymous):

A polygon has n sides. Two of its angles are right angles. Each of the remaining angles equal to 150 degrees. calculate n. Please help! =)

jimthompson5910 (jim_thompson5910):

Two of its angles are right angles...so we have 2*90 = 180 degrees so far

jimthompson5910 (jim_thompson5910):

The remaining angles are 150 degrees there are n-2 angles left since we used up 2 angles (above), so the remaining angles add up to 150(n-2)

jimthompson5910 (jim_thompson5910):

in total, all n angles add to 180 + 150(n-2)

jimthompson5910 (jim_thompson5910):

now, you can use the formula S = 180(n-2) to find the sum of the interior angles of any polygon with n sides

jimthompson5910 (jim_thompson5910):

so if all angles add to 180 + 150(n-2) AND they add to 180(n-2) then 180(n-2) = 180 + 150(n-2)

jimthompson5910 (jim_thompson5910):

solve for n

OpenStudy (anonymous):

the answer would be 8 right because i found this online: For a polygon of n sides, the sum of all angles is 180n – 360 degrees We are given that two sides are 90 degrees and the remainder are 150 degrees Therefore 180n – 360 = 150(n – 2) + 90 x 2 = 150n – 300 + 180 180n = 150n + 360 + 180 – 300 30n = 240 n = 8 Therefore it is an octagon with six angles 150 degrees and two right angles

jimthompson5910 (jim_thompson5910):

that is correct

jimthompson5910 (jim_thompson5910):

solving 180(n-2) = 180 + 150(n-2) for n gives you n = 8

OpenStudy (anonymous):

but i am not understanding that procedure =/

jimthompson5910 (jim_thompson5910):

which one: mine or the thing you found online?

OpenStudy (anonymous):

both =/

jimthompson5910 (jim_thompson5910):

ok where are you stuck on mine?

OpenStudy (anonymous):

how did you get the answer 8

jimthompson5910 (jim_thompson5910):

do you know how to solve 180(n-2) = 180 + 150(n-2) for n?

OpenStudy (anonymous):

nope =/

jimthompson5910 (jim_thompson5910):

ok first distribute 180(n-2) = 180 + 150(n-2) 180n-360 = 180 + 150n-300 what's next?

OpenStudy (anonymous):

where did the 300 come from =/

jimthompson5910 (jim_thompson5910):

150 times 2 = 300

jimthompson5910 (jim_thompson5910):

150 times -2 = -300

OpenStudy (anonymous):

ohh

jimthompson5910 (jim_thompson5910):

so what's next?

OpenStudy (anonymous):

you would have to group the numbers =/

jimthompson5910 (jim_thompson5910):

180 - 300 = ???

OpenStudy (anonymous):

-120

jimthompson5910 (jim_thompson5910):

so 180n-360 = 180 + 150n-300 turns into 180n-360 = 150n-120

jimthompson5910 (jim_thompson5910):

then what?

OpenStudy (anonymous):

dont you group them like: 180n-150n=-120+360

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

basically you get the like terms to the same side

jimthompson5910 (jim_thompson5910):

180n - 150n = ???

OpenStudy (anonymous):

30n

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

-120 + 360 = ?

OpenStudy (anonymous):

then the other answer is 240

jimthompson5910 (jim_thompson5910):

so we can say 30n = 240

OpenStudy (anonymous):

then you divide 240 by 30

jimthompson5910 (jim_thompson5910):

which in english translates to 30 times some number equals 240

jimthompson5910 (jim_thompson5910):

perfect

OpenStudy (anonymous):

Thank you very much! Finally i understand!

jimthompson5910 (jim_thompson5910):

ok that's great that you do